OpenStudy (anonymous):

A monkey in a perch 20m high in a tree drops coconut above the head of a zoo keeper as he runs with a speed of 1.5m/s beneath the tree how far behind him in meters does the coconut hit the ground? please help me

OpenStudy (anonymous):

wat is the speed at which the coconut is falling?

OpenStudy (anonymous):

\[20-(9.8/2)t^2=0\] \[\sqrt{20/4.9}\] \[\sqrt{20/4.9}*1.5=3.03m \]

OpenStudy (stormfire1):

This may help you as well...for constant acceleration (ignoring drag etc): \[distance = 1/2 g t ^2\] So you can simply plug in values to get this equation: \[20m = 1/2(9.8 m/s^2)t^2\]Isolate t \[(20m * 2)/9.8m/s^2=t^2\]Take the square of both sides to solve for t which is how long it will take for the coconut to hit the ground (2.02 seconds). From there, you simply figure out where the monkey will be in 2.02 seconds: \[2.02s *1.5m/s =3.03m\]